OpenStudy (anonymous):

A GARDEN HAS THE SHAPE OF AN ISOSCELES RIGHT TRIANGLE THE LENGTH OF THE HYPOTENUSE IS 24 FEET WHAT US THE AREA OF THE GARDEN ?

OpenStudy (anonymous):

go to math group lol this is history

OpenStudy (anonymous):

@Notamathgenius @SWAG @SwaggKiller @marrero @shaik0124

OpenStudy (swag):

Hi, I can see that you are a new member here at OpenStudy so, I would first like to say Welcome to OpenStudy, I would like to point you to the chat pods these chat pods are where you can make new friends and talk to new people just like you, I would also like to emphasize our "NON-CHEATING POLICY" we ask here at OpenStudy that you do NOT post exam/Test questions if you are caught doing this we will notify your current school I would also like to ask you to check out the OpenStudy "Code of Conduct" http://openstudy.com/code-of-conduct Please read this carefully and thoroughly. If you need any help feel free to message me or any other ambassador. Welcome to the OpenStudy Community! Ambassador SWAG

OpenStudy (anonymous):

and barbie: u have to get used to swag.

OpenStudy (anonymous):

u learn to.

OpenStudy (anonymous):

I'm not cheating and wow

OpenStudy (anonymous):

@swag what is the role of an embassador??

OpenStudy (anonymous):

to help

OpenStudy (anonymous):

and he wasnt, u have to read the terms of service and such be4 u create an OS.

OpenStudy (swag):

This a welcome not a warning.

OpenStudy (swag):

is the only part u read the middle? geesh

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

I think barbie just got overwhelmed

OpenStudy (swag):

The Ambassadors Greet New Users & Make sure they follow CoC

OpenStudy (anonymous):

swag how old r u?

OpenStudy (anonymous):

Oh sweet lol i couldn't ask you in a message because it wouldn't let me send you one. said something like you only accept messages from people you fanned

OpenStudy (anonymous):

so do we help ?

OpenStudy (anonymous):

@barbie15 as in?

OpenStudy (anonymous):

@redred413

OpenStudy (anonymous):

How old are y'all ?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

19

OpenStudy (anonymous):

i need some help on any of thses questions i have please

OpenStudy (anonymous):

and im 16!!!

OpenStudy (anonymous):

this wasn't closed??

OpenStudy (anonymous):

24^2=a^2+b^2 576=a^2+b^2 576=288+288 sqrt 288=16.97 a=1/2LW a=.5*16.97*16.97 a=144 sqft

OpenStudy (anonymous):

thnx